6, what is the value of 15-year corporate bonds, with a coupon rate of 9%, if current interest rates on similar bonds is 8%? How much would the value change if interest rates increased to 10%? Under what conditions will this bond trade at par (face value)?
Answer #1

Assumptions :

Bond Face value = $1,000 and coupons are paid semi-annually.

Coupon payment will be1,000 * 9% / 2 = $45

Interest rate is 8% but since we are assuming semi annual coupon payment, it will be 4% and period will be 30 (15 year * 2).

Value of Bond will be

45*(1-(1.04^-30))/0.04 + 1000/(1.04^30) = $1,086.46

If the market rate increases to 10% :

45*(1-(1.05^-30))/0.05 + 1000/(1.05^30) = $923.14

The bond will trade at par if the market interest rate is equal to coupon rate. So in this case when market interest rate becomes 9%, the bond will trade at par.